If four out of five doctors prefer fast food, than at a convention of 455 doctors, how many probably ate fast food for lunch? Set up an equation and solve.
thanks thanks

Answer by stanbon(75887)

If four out of five doctors prefer fast food, than at a convention of 455 doctors, how many probably ate fast food for lunch? Set up an equation and solve.
--------------
Let x = # preferring fast food
-----
Equation:
x = (4/5)455 = 4*91 = 364
=============================
Cheers,
Stan H.